# Explainable Concept Generation through Vision-Language Preference Learning

### v2

@@ 4.5 Applications and generalizability
To demonstrate the generalizability of the proposed algorithm, we extended RLPO to generate words in sentiment analysis in NLP. We made use of Mistral-7B Instruct model to generate synonyms of seed prompts and optimized the language model based on preferences from TextCNN model pre-trained on IMDB sentiment dataset. Fig. 10 highlights relevant words with their importance score in the input. More details about the experiment is present in Appendix D.4.
@@ 5 Limitations and Conclusions
The process of navigating an infinitely large concept space and generating explainable concepts from textual inputs presents several challenges, particularly when dealing with complex, high-level concepts. We showed how deep RL can guide SD efficiently to navigate this space. However, RLPO also suffers from several limitations. First, this analysis cannot be performed real-time since generating images from SD, learning the DQN, and fine-tuning the SD with preferences takes some time. Also, the concepts that our algorithm generates can be diverse as it tries to reveal the concepts inherent to the NUT, making it less domain-specific (e.g., for a medical application, there is a chance it might generate non-medical images if the NUT activations get excited for non-medical data). As a future extension, we hope to input preferences from both TCAV and application experts while optimizing, making generated explanations even more aligned to specific applications. Despite the challenges, our results show how to leverage the strengths of visual representations and adaptive learning to provide intuitive and effective solutions for understanding complex, high-level concepts in black-box neural networks.
@@
Appendix A Related Work
Use of VLM in explaination: Recent advancements in Vision-Language models (VLMs) have open the doors for the use of VLMs in multiple domains, mainly because of their ability to generalize over large amount of data, they can be leveraged to obtain useful information. Work by Sun et al. (2024) present a novel method combining the Segment Anything Model (SAM) with concept-based explanations, called Explain Any Concept (EAC). This method uses SAM for precise instance segmentation to automatically extract concept sets from images, then it employs a lightweight surrogate model to efficiently explain decision made by any neural network based on extracted concepts. Another work by Yan et al. Yan et al. (2023) introduced Learning Concise and Descriptive Attributes (LCDA), which leverages Large Language Models (LLMs) to query a set of attributes describing each class and then use that information with vision-language models to classify images. They highlight in their paper that with a concise set of attributes, they can improve the classifier’s performance and also increase interpretability and interactivity for end user.
Appendix B Definition, Theorems, and Proves
Comparative Overlap of Human-Interpretable and Generative Model Concepts in Neural Understanding Tasks (NUT). We formalize this with reference to Fig. 2.
@@ Appendix B Definition, Theorems, and Proves
Let the set of human-interpretable concepts that the NUT has learned be $\mathcal{C}_{N}$, and the concept sets human collected, retrieved though segmentation, and generated using a generative model be $\mathcal{C}_{H}$, $\mathcal{C}_{R}$, and $\mathcal{C}_{G}$, respectively. Then, $|\mathcal{C}_{G}\cap\mathcal{C}_{N}|\geq|\mathcal{C}_{H}\cap\mathcal{C}_{N}|\geq 0$ and $|\mathcal{C}_{G}\cap\mathcal{C}_{N}|\geq|\mathcal{C}_{R}\cap\mathcal{C}_{N}|\geq 0$.
@@
where $\mathcal{C}_{N}$ represents the set of concepts learned by the NUT.
@@
The Effects of DQN-DPO-based Concept Space Traversal. We now formalize what concepts the DQN has learned, with reference to Fig. 1.
@@
As $\xi$ increases, the reward function proportionally amplifies, particularly enhancing the significance of outcomes near $t_{\eta}$.
@@
As we go closer to the concept class, $|\mathcal{C}_{G}\cap\mathcal{C}_{N}|$ becomes larger for generated concepts $\mathcal{C}_{G}$ and NUT’s internal concepts, $\mathcal{C}_{N}$.
@@ B.1 Definitions
Average Normalized Count (ANC): The ANC is a measure of the central tendency of the normalized action frequencies within a distribution. It provides insight into how the actions are distributed relative to the overall frequency distribution. A high ANC indicates that, on average, the action frequencies are relatively large, meaning that certain actions are more dominant. Conversely, a low ANC suggests that the actions are more evenly distributed or that lower-frequency actions have a more significant presence. Given by $\frac{1}{n}\sum_{i=1}^{n}\frac{f_{i}}{\sum_{j=1}^{n}f_{j}}$, where $f_{i}$ is the frequency of action $i$.
@@
Appendix C Methodology
C.1 Machine learning models we use
Neural Network Under Test ($f(.)$) : We test RLPO for all the different classification models given below.
@@ C.1 Machine learning models we use
    ResNet50: We utilized a pretrained model from PyTorch torchvision pretrained models with weights initialized from ResNet50_Weights.IMAGENET1K_V2.
@@
TCAV logistic model : We utilized a logistic regression model to address classification tasks in TCAV instead of the default SGD (Stochastic Gradient Descent) classifier. This decision was based on our observation that the SGD classifier produced high variance TCAV (Testing with Concept Activation Vectors) scores, which indicated inconsistent model behavior across different runs. We configured the model to perform a maximum of 1000 iterations (max_iter=1000).
Stable Diffusion v1-5 with LoRA : We used our base generation model as SD v1-5 and updated its weights using LoRA during preference optimization step. This version of SD was finetuned from SD v1-2 on "laion-aesthetics v2 5+" dataset with 10% drop in text-conditioning for better CFG sampling. In our experiments, we kept LoRA rank to 8 with a scaling factor of 8 and initial weights were defined from a gaussian distribution. We only targeted the transformer modules of U-Net in the SD architecture.
@@
C.2 RLPO algorithm
Algorithm 2 shows the complete algorithm of the algorithm shown in Section 3 algorithm 1.
@@ C.2 RLPO algorithm
1:Input: Set of test images, $f(.)$
2:Initialize Q-network $Q_{\theta}(s,a)$ with random weights $\theta$
3:Initialize replay buffer $\mathcal{D}$ and adaptive parameter $\xi\leftarrow 0.1$
4:for each episode do
5:     for each time step $t$ do
6:         Observe state $s_{t}$ and select action $a_{t}$ based on $Q$ ($\epsilon$-greedy)
7:         Execute $a_{t}$ and generate 10 images, divided into two groups $G_{1}$ and $G_{2}$
8:         Evaluate TCAV scores $\overline{TCAV}_{1}$ and $\overline{TCAV}_{2}$
9:         if  $\max(\overline{TCAV}_{1},\overline{TCAV}_{2})\leq 0.7$  then
10:              Update policy to favor higher $\overline{TCAV}$ group and perform DPO
11:              Update $\xi\leftarrow\min(1,\xi+\text{increment})$
12:         else
13:              Set $\xi\leftarrow 1$
14:         end if
15:         Compute reward $r_{t}=\xi\cdot\max(\overline{TCAV}_{1},\overline{TCAV}_{2})$
16:         Store transition $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{D}$
17:         Sample a mini-batch from $\mathcal{D}$
18:         for each sampled transition $(s_{i},a_{i},r_{i},s_{i+1})$ do
19:              Compute target $y_{i}=r_{i}+\gamma\max_{a^{\prime}}Q_{\theta^{\prime}}(s_{i+1},a^{\prime})$
20:         end for
21:         Compute loss $L(\theta)=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-Q_{\theta}(s_{i},a_{i}))^{2}$
22:         Perform a gradient descent step to update $\theta$
23:         Periodically update target network: $\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime}$
24:     end for
25:end for
26:Output: Set of concept images
C.3 Preprocessing for generating the action space
@@
Each patch from the test images is passed to the VQA model to extract relevant and useful information about the corresponding class. In this study, we choose BLIP Li et al. (2022) as our VAQ model. We posed a set of targeted questions to the VQA model, aiming to gain insights into the class-specific features represented in the patches. The questions are designed to probe various aspects of the image patches, helping the model focus on class-defining attributes.
@@
We then remove stop words and duplicates from the generated responses using lemmantizing and perform a cross-similarity check using CLIP between all the unique words and further filtered words which are more than 95% similar. To further select most relevant keywords to the class images, we perform a VLM check using class images and the extracted keyword to get the softmax score of how much the keyword and image are related. This score is then averaged over all the class images and this average is use to sort the keywords. Now, from the sorted keywords, we select top 20 keywords as our RL action space. The cross-similarity and VLM check are inspired from Zang et al. (2024) where they used a similar filtering setup to remove potentially useless concepts.
C.4 Exploration of Random Gibberish Prompts as Seed Prompts
In this experiment we don’t use a VQA to get seed prompts. We choose a random list of incoherent prompts, example shown in Fig. 12. We found that for these prompts it takes a really long time to get some meaningful explanation and in most cases lead to random generation, thus showing importance of starting from proxy concepts.
@@ C.4 Exploration of Random Gibberish Prompts as Seed Prompts
C.5 Preference optimization update for state space
@@
The candidate concepts serve as the initial states for the RL agent. From these initial states, the agent takes actions $a\in\text{Keywords}$ that leads to multiple subsequent possible states using $g(.)$. These states are then grouped, and the group’s sensitivity is compared against INUT using TCAV scores. A higher TCAV score suggests higher sensitivity, indicating that the group is more aligned with INUT.
We employ preference optimization over the grouped states to guide states towards INUT. To prevent the model from skipping over explainable states and directly reaching INUT, we introduce a threshold that limits the application of preference optimization at each step as shown in equation 17.
@@
To optimize $g(.)$ to find better proxies, for each step in the environment we utilized average TCAV scores $\overline{TCAV}_{1}$ and $\overline{TCAV}_{2}$ from $G_{1}$ and $G_{2}$ to decide between preferred and unpreferred concepts. Lets say $\overline{TCAV}_{1}\succ\overline{TCAV}_{2}$, than we optimize $g(.)$ over the sample S defined as $S=\{(a,x^{g1}_{0},x^{g2}_{0})\}$, where $x^{g1}_{0}$ and $x^{g2}_{0}$ are the sample points from the groups on action $a$. We optimize $g(.)$ using objective 18 to get a new optimzed $g^{\prime}(.)$  Wallace et al. (2023).
@@
C.6 TCAV setting for different NUT models
We tested NUT models on different layers and classes and the summary of our setting across different NUT models is described in table 5.
Table 5: TCAV setting across different NUT models
@@ C.6 TCAV setting for different NUT models
ResNet50 penultimate layer Jay, Tiger, Rabbit & Zebra
@@
Appendix D Experiments
D.1 Computing resources
@@
D.2 Human and LLM-based AI Feedback Mechanisms
@@
AI feedback: GPT-4 is leveraged to evaluate the explanatory power of image sets by focusing on concepts related to a target class, a method that aligns with the growing trend of incorporating AI-driven feedback Bai et al. (2022). This approach involves sending a structured prompt to an LLM, asking it to score how well two sets of images explain a target class using a specified concept. The process involves the following.
@@
Human feedback: In this experiment, 7 computer science majors provided live feedback after each step of a reinforcement learning process, leveraging their prior knowledge of reinforcement learning with human feedback (RLHF) mechanisms Christiano et al. (2017). The feedback from all participants was averaged to serve as the reward for each step in the RL process. Given the abstract nature of the initial concepts, participants needed to take time to thoughtfully assess each step, which contributed to a lengthier feedback cycle.
D.3 Additional results and analysis
To validate our method for its ability to generate concepts, we tested it with different models and classes. We started it on traditional models, GoogleNet and InceptionV3, and then extended it to transformer-based models, Vision Transformer (ViT) and Swin Transformer, pre-trained on ILSRVC2012 data set (ImageNet) Krizhevsky et al. (2017). We show additional plot in various classes shown in Fig 13,14,15,16,17.

### v3

@@ 4.7 Can RLPO Generalized Beyond Images?
To demonstrate the generalizability of the proposed algorithm, we extended RLPO to generate words to explain sentiment analysis in NLP. We made use of Mistral-7B Instruct model to generate synonyms of seed prompts and optimized the language model based on preferences from TextCNN model pre-trained on IMDB sentiment dataset. Fig. 9 highlights relevant words with their importance score in the input. More details are provided in Appendix C.7.
Figure 9: We use sentiment analysis in NLP to show the generalizability of RLPO. Concepts tend to be synonyms and the numbers indicate the TCAV scores.
@@ 5 Limitations and Conclusions
In this work, we introduced Reinforcement Learning-based Preference Optimization (RLPO) to automatically articulate concepts that explain the internal representations of neural networks. We demonstrated how RLPO can guide a vision-language generative model to navigate an infinitely large concept space, addressing the challenge of manually curating concept image sets. However, our approach also suffers from some limitations. Firstly, while RLPO is designed to iteratively refine these prompts and can drift towards more model-relevant abstractions beyond the original scope, the quality and nature of the initial seeds can influence the search trajectory and efficiency. Increasing the number of seeds helps generate better concepts. Additionally, the quality and diversity of the generated concepts are also influenced by the generative model being used. The pre-training data of such models can introduce biases, and their capacity to accurately render certain concepts. While this dependency is not unique to generative approach, retrieval-based methods are similarly constrained by available data and human cognitive biases, explicitly addressing and potentially mitigating these inherited biases in the context of XAI is an important direction.
The concepts that our algorithm generates can be diverse as it tries to reveal the concepts inherent to the $f(\cdot)$, making it less domain-specific (e.g., for a medical application, there is a chance it might generate non-medical images if the $f(\cdot)$ activations get excited for non-medical data). As a future extension, we aim to input preferences from both TCAV and domain experts while optimizing, making generated explanations even more aligned to specific applications. Despite these challenges, our results demonstrate how to leverage the strengths of visual representations and adaptive learning to provide intuitive and effective solutions for understanding complex, high-level concepts in neural networks.
Impact Statement
Our paper is a generic algorithmic contribution, and we do not foresee direct negative societal impacts. As a positive trait, our approach can potentially be used to understand and explain model bias.
@@
In this appendix, we describe the theorems, algorithm, metrics, experimental setup and additional results.
Appendix A Definition, Theorems, and Proves
Comparative Overlap of Human-Interpretable and Generative Model Concepts in Neural Understanding Tasks ($f(\cdot)$). We formalize this with reference to Fig. 1.
@@ Appendix A Definition, Theorems, and Proves
Let the set of human-interpretable concepts that the $f(\cdot)$ has learned be $\mathcal{C}_{N}$, and the concept sets human collected, retrieved though segmentation, and generated using a generative model be $\mathcal{C}_{H}$, $\mathcal{C}_{R}$, and $\mathcal{C}_{G}$, respectively. Then, $|\mathcal{C}_{G}\cap\mathcal{C}_{N}|\geq|\mathcal{C}_{H}\cap\mathcal{C}_{N}|\geq 0$ and $|\mathcal{C}_{G}\cap\mathcal{C}_{N}|\geq|\mathcal{C}_{R}\cap\mathcal{C}_{N}|\geq 0$.
@@
where $\mathcal{C}_{N}$ represents the set of concepts learned by $f(\cdot)$.
@@
The Effects of DQN-PO based Concept Space Traversal. We now formalize what concepts the DQN has learned, with reference to Fig. 2.
@@
As $\xi$ increases, the reward function proportionally amplifies, particularly enhancing the significance of outcomes near $t_{\eta}$, which marks the point beyond which TCAV scores are always 1.
@@
As we go closer to the concept class, $|\mathcal{C}_{G}\cap\mathcal{C}_{N}|$ becomes larger for generated concepts $\mathcal{C}_{G}$ and $f(\cdot)$’s internal concepts, $\mathcal{C}_{N}$.
@@ A.1 Definitions
Average Normalized Count (ANC): The ANC is a measure of the central tendency of the normalized action frequencies within a distribution. It provides insight into how the actions are distributed relative to the overall frequency distribution. A high ANC indicates that, on average, the action frequencies are relatively large, meaning that certain actions are more dominant. Conversely, a low ANC suggests that the actions are low and only a few high frequent actions are present. Given by $\frac{1}{n\cdot\max(f)}\sum_{i=1}^{n}f_{i}$, where $f_{i}$ is the frequency of action $i$.
@@
Appendix B Methodology
B.1 Machine Learning Models We Use
Neural Network Under Test ($f(\cdot)$): We test RLPO for all the different classification models given below.
@@ B.1 Machine Learning Models We Use
TCAV Logistic Model: We utilized a logistic regression model to address classification tasks in TCAV instead of the default SGD (Stochastic Gradient Descent) classifier. This decision was based on our observation that the SGD classifier produced high variance TCAV (Testing with Concept Activation Vectors) scores, which indicated inconsistent model behavior across different runs. As demonstrated in Table 7, the standard deviations for SGD-derived scores are frequently larger than those obtained with the logistic model, underscoring the latter’s improved stability. We configured the model to perform a maximum of 1000 iterations (max_iter=1000).
Table 7: TCAV Scores (Concept/Random) for Different Models, Layers, and Classifiers. Scores are presented as mean $\pm$ standard deviation across 5 runs with random seed. Concepts used in these experiments are from the ones provided by the authors of TCAV paper.
Model Layer Stripes/Random TCAV Score Dots/Random TCAV Score
SGD Logistic SGD Logistic
GoogleNet inception3a $0.662\pm 0.03$ / $0.338\pm 0.03$ $0.67\pm 0.00$ / $0.33\pm 0.00$ $0.36\pm 0.05$ / $0.64\pm 0.05$ $0.33\pm 0.00$ / $0.67\pm 0.00$
inception4e $0.992\pm 0.01$ / $0.008\pm 0.01$ $1.00\pm 0.00$ / $00.00\pm 0$ 0.00 $0.01\pm 0.007$ / $0.99\pm 0.007$ $00.00\pm 0$ 0.00 / $1.00\pm 0.00$
ResNet50 layer3 $0.796\pm 0.02$ / $0.204\pm 0.02$ $0.78\pm 0.00$ / $0.22\pm 0.00$ $0.078\pm 0.07$ / $0.922\pm 0.07$ $00.00\pm 0$ 0.00 / $1.00\pm 0.00$
layer4 $1.000\pm 0.00$ / $00.000\pm 0$ 0.00 $1.00\pm 0.00$ / $00.00\pm 0$ 0.00 $0.60\pm 0.54$ / $0.40\pm 0.54$ $00.00\pm 0$ 0.00 / $1.00\pm 0.00$
Stable Diffusion v1-5 with LoRA: We used our base generation model as SD v1-5 and updated its weights using LoRA during preference optimization step. This version of SD was finetuned from SD v1-2 on “laion-aesthetics v2 5+” dataset with 10% drop in text-conditioning for better CFG sampling. In our experiments, we kept LoRA rank to 8 with a scaling factor of 8 and initial weights were defined from a gaussian distribution. We only targeted the transformer modules of U-Net in the SD architecture.
@@
B.2 RLPO Algorithm
Algorithm 2 presents the detailed version of the algorithm introduced in Section 3, Algorithm 1.
@@ B.2 RLPO Algorithm
  Input: Set of test images, $f(\cdot)$
  Initialize Q-network $Q_{\theta}(s,a)$ with random weights $\theta$
  Initialize replay buffer $\mathcal{D}$ and adaptive parameter $\xi\leftarrow 0.1$
  for each episode do
     for each time step $t$ do
        Observe state $s_{t}$ and select action $a_{t}$ based on $Q$ ($\epsilon$-greedy)
        Execute $a_{t}$ and generate 10 images, divided into two groups $G_{1}$ and $G_{2}$
        Evaluate TCAV scores $\overline{TCAV}_{1}$ and $\overline{TCAV}_{2}$
        if $\max(\overline{TCAV}_{1},\overline{TCAV}_{2})\leq 0.7$ then
           Update policy to favor higher $\overline{TCAV}$ group and perform DPO
           Update $\xi\leftarrow\min(1,\xi+\text{increment})$
        else
           Set $\xi\leftarrow 1$
        end if
        Compute reward $r_{t}=\xi\cdot\max(\overline{TCAV}_{1},\overline{TCAV}_{2})$
        Store transition $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{D}$
        Sample a mini-batch from $\mathcal{D}$
        for each sampled transition $(s_{i},a_{i},r_{i},s_{i+1})$ do
           Compute target $y_{i}=r_{i}+\gamma\max_{a^{\prime}}Q_{\theta^{\prime}}(s_{i+1},a^{\prime})$
        end for
        Compute loss $L(\theta)=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-Q_{\theta}(s_{i},a_{i}))^{2}$
        Perform a gradient descent step to update $\theta$
        Periodically update target network: $\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime}$
     end for
  end for
  Output: Set of concept images
B.3 Preprocessing for Generating the Action Space
@@
Each patch from the test images is passed to the VQA model to extract relevant and useful information about the corresponding class. In this study, we choose BLIP (Li et al., 2022) as our VAQ model. We posed a set of targeted questions to the VQA model, aiming to gain insights into the class-specific features represented in the patches. The questions are designed to probe various aspects of the image patches, helping the model focus on class-defining attributes.
@@
We then remove stop words and duplicates from the generated responses using lemmantizing and perform a cross-similarity check using CLIP between all the unique words and further filtered words which are more than 95% similar. To further select most relevant keywords to the class images, we perform a VLM check using class images and the extracted keyword to get the softmax score of how much the keyword and image are related. This score is then averaged over all the class images and this average is use to sort the keywords. Now, from the sorted keywords, we select top 20 keywords as our RL action space. The cross-similarity and VLM check are inspired from (Zang et al., 2025) where they used a similar filtering setup to remove potentially useless concepts.
B.4 Exploration of Random Gibberish Prompts as Seed Prompts
In this experiment we don’t use a VQA to get seed prompts. We choose a random list of incoherent prompts, example shown in Fig. 11. We found that for these prompts it takes a really long time to get some meaningful explanation and in most cases lead to random generation, thus showing importance of starting from good proxy concepts.
@@ B.4 Exploration of Random Gibberish Prompts as Seed Prompts
B.5 Preference Optimization Update for State Space
@@
The candidate concepts serve as the initial states for the RL agent. From these initial states, the agent takes actions $a\in\text{Keywords}$ that leads to multiple subsequent possible states using $g(.)$. These states are then grouped, and the group’s sensitivity is compared against Inputs of $f(\cdot)$ using TCAV scores. A higher TCAV score suggests higher sensitivity, indicating that the group is more aligned with $f(\cdot)$’s inputs.
We employ preference optimization over the grouped states to guide states towards explainable concepts. To prevent the model from skipping over explainable states and directly reaching the input domain, we introduce a threshold that limits the application of preference optimization at each step as shown in equation 17.
@@
To optimize $g(.)$ to find better proxies, for each step in the environment we utilized average TCAV scores $\overline{TCAV}_{1}$ and $\overline{TCAV}_{2}$ from $G_{1}$ and $G_{2}$ to decide between preferred and unpreferred concepts. Lets say $\overline{TCAV}_{1}\succ\overline{TCAV}_{2}$, than we optimize $g(.)$ over the sample S defined as $S=\{(a,x^{g1}_{0},x^{g2}_{0})\}$, where $x^{g1}_{0}$ and $x^{g2}_{0}$ are the sample points from the groups on action $a$. We optimize $g(.)$ using objective 18 to get a new optimzed $g^{\prime}(.)$  (Wallace et al., 2024).
@@
B.6 TCAV Setting for Different Models
We tested different models on different layers and classes and the summary of our setting across different models is described in table 8.
Table 8: TCAV setting across different models
@@ B.6 TCAV Setting for Different Models
Appendix C Experiments
C.1 Computing Resources
@@
C.2 Human and LLM-Based AI Feedback Mechanisms
@@
AI Feedback: GPT-4 is leveraged to evaluate the explanatory power of image sets by focusing on concepts related to a target class, a method that aligns with the growing trend of incorporating AI-driven feedback (Bai et al., 2022). This approach involves sending a structured prompt to an LLM, asking it to score how well two sets of images explain a target class using a specified concept. The process involves the following.
@@
Human Feedback: In this experiment, eight computer science majors provided live feedback after each step of a reinforcement learning process, leveraging their prior knowledge of reinforcement learning with human feedback (RLHF) mechanisms (Christiano et al., 2017). The feedback from all participants was averaged to serve as the reward for each step in the RL process. Given the abstract nature of the initial concepts, participants needed to take time to thoughtfully assess each step, which contributed to a lengthier feedback cycle.
C.3 Additional Results and Analysis
To validate our method for its ability to generate concepts, we tested it with different models and classes. We started it on traditional models, GoogleNet and InceptionV3, and then extended it to transformer-based models, Vision Transformer (ViT) and Swin Transformer, pre-trained on ILSRVC2012 data set (ImageNet) (Krizhevsky et al., 2017). We show additional plot in various classes shown in Fig 12,13,14,15,16.